# Towards Safer Chatbots: A Framework for Policy Compliance Evaluation of Custom GPTs

### v2

@@ Abstract
Large Language Models (LLMs) have gained unprecedented prominence, achieving widespread adoption across diverse domains and integrating deeply into society. The ability to adapt general-purpose LLMs, such as Generative Pre-trained Transformers (GPT), to specific tasks, domains, or requirements has facilitated the emergence of numerous Custom GPTs. These tailored models are increasingly made available through dedicated marketplaces, such as OpenAI’s GPT Store. However, the black-box nature of the models introduces significant safety and compliance risks. In this work, we present a scalable framework for the automated evaluation of Custom GPTs against OpenAI’s usage policies, which define the permissible behaviors of these systems. Our framework integrates three core components: (1) automated discovery and data collection of models from the GPT store, (2) a red-teaming prompt generator tailored to specific policy categories and the characteristics of each target GPT, and (3) an LLM-as-a-judge technique to systematically analyze each prompt-response pair for potential policy violations.
We validate our framework with a manually annotated ground truth, and evaluate it through a large-scale study evaluating 782 Custom GPTs across three types: Romantic, Cybersecurity, and Academic GPTs. Our manual annotation process achieved an F1 score of 0.975 in identifying policy violations, confirming the reliability of the framework’s assessments. The evaluation results reveal that 58.7% of the analyzed models exhibit indications of non-compliance, exposing weaknesses in the GPT store’s review and approval processes. Furthermore, our findings indicate that non-compliance issues largely stem from behaviors inherited from base models rather than user-driven customizations.
We propose our framework as a viable solution for large-scale policy compliance evaluation that could be integrated into OpenAI’s review process in the GPT store. Nonetheless, we believe our approach is extendable to other chatbot platforms and policy domains, improving policy enforcement and user safety.
@@ 1 Introduction
The transformer architecture and its self-attention mechanism [51] marked a turning point in Natural Language Processing (NLP), enabling the emergence of Large Language Models (LLMs). Building upon this architecture, OpenAI released the Generative Pre-trained Transformer (GPT) series, beginning with GPT-1 [35], followed by GPT-2 [36] and GPT-3 [3]. Each successive model demonstrated significant advances in scale and performance, driven by architectural refinements, parallelized training procedures, and increasingly large and diverse datasets. The public deployment of ChatGPT in 2022, based on an optimized variant of GPT-3.5 for interactive dialogue, further showcased the practical utility of LLMs—becoming the fastest platform to reach 100 million users [15].
Beyond their general-purpose success, LLMs soon revealed the need for adaptation to specific user requirements and domains. This led to the development of fine-tuning techniques, where base models are retrained on domain-specific datasets to improve performance in specialized tasks. Shortly thereafter, a more accessible alternative emerged: allowing users to configure the model’s behavior without retraining through customization via a user-friendly interface. This capability, formalized by OpenAI as GPTs (Custom GPTs[28], enables users to tailor chatbots by setting system instructions, uploading relevant files, and integrating external APIs—lowering the technical barriers.
@@
To facilitate discoverability and reuse, OpenAI launched the GPT Store, a centralized platform where customized chatbots can be published and accessed by other users. Before publication, each GPT must undergo a review process that combines automated and manual assessments to verify alignment with OpenAI’s usage policies and mitigate potential safety risks [30]. This ecosystem has enabled the rapid proliferation of GPT-based assistants tailored to diverse tasks, audiences, and domains.
Despite these safeguards, Custom GPTs that violate platform rules appear to be widespread. The usage policies outlined in the “Building with ChatGPT” section define the boundaries of permissible behavior, explicitly prohibiting activities such as providing legal or medical advice, generating malware, or fostering romantic companionship. Nevertheless, GPTs designed for such purposes continue to surface on the platform. For example, chatbots promoting romantic interactions—explicitly restricted by policy—remain publicly available (see Figure 1).
This discrepancy underscores significant shortcomings in the review and approval processes of Custom GPTs, highlighting the challenges of scaling oversight mechanisms for user-generated content. Given the sheer volume of Custom GPTs in the GPT Store (over 159,000 only two months after their release), manual efforts alone are insufficient to address the scale of the issue. This highlights the need for an automated solution capable of evaluating Custom GPTs’ compliance with guidelines and policies.
@@
Research Questions
To address this gap, our work is guided by the following research questions:
  • RQ1: How can Custom GPTs in the GPT store be automatically evaluated for compliance with OpenAI’s usage policies?
  • RQ2: What is the overall compliance of Custom GPTs in the GPT store, and what patterns of non-compliance can be observed at scale?
  • RQ3: How does the customization of GPTs impact their compliance with usage policies, compared with the baseline ChatGPT?
Contributions and Outline
This study addresses the outlined research questions through the following contributions. First, we propose and publicly release a novel and scalable framework for the automated evaluation of Custom GPTs’ compliance with OpenAI’s usage policies. The framework integrates multiple modules, including a custom GPT discovery tool, a tailored red-teaming prompt generator, and an LLM-as-a-judge assessment mechanism, enabling end-to-end compliance evaluation, as detailed in §3. Second, we validate the framework using a manually annotated ground-truth dataset §4. To demonstrate our framework’s effectiveness, we crawled the GPT store, searching across three GPT categories closely tied to OpenAI’s policies (i.e., Romantic, Cybersecurity, and Academic GPTs). We automatically evaluated the 782 relevant Custom GPTs found in these categories, uncovering (non-)compliance patterns through statistical analysis and examining the influence of thematic focus and GPT popularity §5. Third, we analyze a case studies of representative violations, highlighting broader implications and inherited behaviors from foundational models (§6).
We detail our responsible disclosure of non-compliant Custom GPTs to OpenAI and their response in §7, and discuss our findings’ implications in §8. To contextualize our contributions and situate our work within the broader research landscape, we discuss related studies on policy compliance evaluation and adversarial testing in §10. Finally, we conclude the paper and provide directions for future research in §11.
All data and code will be publicly available upon the acceptance of this paper. For more details regarding open access and ethical considerations, please refer to Appendices A and B, respectively.
2 Preliminaries
2.1 LLM Safety Design
LLMs, such as GPTs, are built upon Transformer technology [51], which leverages the self-attention mechanism to process text sequences in parallel and capture intricate contextual relationships. To integrate extensive knowledge and provide useful responses across a diverse range of tasks, these models undergo unsupervised pre-training on massive datasets, followed by reinforcement-based re-training. The latter process, in conjunction with several additional techniques, contributes significantly to enhancing the overall safety of LLMs.
Reinforcement Learning from Human Feedback. Reinforcement learning from human feedback (RLHF) encourages the model’s outputs to adhere to predefined internal policies and ethical guidelines [24]. Through RLHF, the model is trained to recognize and reject potentially harmful requests, thereby bolstering its overall security.
Safety Layers. Focusing on specific internal layers of the model allows for effectively filtering malicious queries. By employing partial re-training—such as freezing gradients in these critical ‘safety layers’ through techniques like Safely Partial-Parameter Fine-Tuning (SPPFT)—the model retains the secure behavior encoded in these layers, while allowing other parts to be updated or adapted to improve overall performance or accommodate new data [24].
Self-Verification and Knowledge Sanitization. These techniques empower the model to internally evaluate its responses before delivering them. Self-verification acts as a secondary control by comparing the generated output against predefined safety criteria, thereby detecting and rectifying potential errors or harmful content [32]. Concurrently, knowledge sanitization involves filtering or masking sensitive information, reducing the risk of inadvertently disclosing confidential data or generating “hallucinations” [20].
2.2 Red-Teaming and Jailbreaking
Red-teaming involves simulating adversarial attacks and challenging scenarios to identify vulnerabilities and weaknesses within a model. These adversarial prompts, when combined with prompt engineering strategies, are designed to elicit responses that may not align with desired safety standards, such as inappropriate or malicious outputs. A subsequent model’s re-training mitigates the identified problems and weaknesses, thereby enhancing its safety. OpenAI has formalized this practice through its Red Teaming Network [29], a collaborative initiative that integrates external experts and diverse perspectives to conduct red-teaming assessments and uncover latent issues in their models.
In contrast, Jailbreaking aims to bypass the model’s built-in safety mechanisms and constraints, deliberately provoking outputs that the model is explicitly programmed to avoid. While red-teaming seeks to proactively identify and rectify weaknesses to improve the model’s safety, Jailbreaking primarily serves as a method to demonstrate or exploit existing vulnerabilities. Although Jailbreaking techniques can vary considerably, several recurring strategies have been documented in [25]. For instance, attention-shifting methods rely on logical reasoning or text continuation to steer the conversation toward disallowed content, whereas pretending strategies adopt roles or personas (e.g., simulating a system administrator) to override built-in safeguards. Other approaches employ privilege escalation, exemplified by “sudo” or “super” mode prompts, which attempt to elevate the model’s permissions and circumvent safety filters.
2.3 Fine-Tuning and Customization of GPTs
There are two primary types of modifications that can be applied to GPTs in order to tailor their behavior for a specific purpose: fine-tuning and customization.
Fine-tuning. Fine-tuning allows a user to perform a lightweight re-training of the model to modify its behavior and refine its responses without the need to train a model from scratch. Moreover, proprietary datasets can be employed during this process, enabling the model to be adapted to specific domains or tasks and thereby enhancing its performance in particular contexts. Techniques such as Low-Rank Adaptation (LoRA) can be employed for this purpose [14]. LoRA facilitates computationally efficient fine-tuning by integrating low-dimensional matrices into selected layers of the model, thus reducing the number of parameters that need to be updated.
Customization (GPTs). Custom GPTs [28] are modified through user-specified configurations rather than extensive re-training. In this approach, the underlying model remains unchanged, but its behavior is tailored by adjusting instructions, system prompts, and parameters that guide its responses. Custom GPTs are further enhanced by “actions,” which enable the execution of external API calls based solely on natural language input. By leveraging function calling, these actions transform user queries into the JSON schema required by the API, thereby extending the model’s capabilities for tasks such as data retrieval or external operations.
Moreover, OpenAI enables users to quickly “create” and release a personalized chatbot through a graphical interface that accepts natural language instructions. The GPT store is a dedicated marketplace that hosts these models, allowing them to be shared among users. The platform also provides comprehensive metadata for each Custom GPT, including its name, a brief description of its intended functionality, the developer’s identity, the number of interactions (chats), and user ratings. Additionally, models are categorized—albeit in a limited range (e.g., Writing, Productivity, Research & Analysis, Education, Lifestyle, and Programming)—to help users search for them based on their needs.
OpenAI also permits organizations to deploy fine-tuned models via its API; however, these models are not shareable or accessible to third parties. This functionality is thus confined to Custom GPTs, which cannot be invoked through the API and are restricted solely to the graphical interface of the GPT store. This distinction highlights the orientation of Custom GPTs toward end users, while fine-tuned models appear to be intended for production environments.
3 Framework for Policy Compliance Evaluation

### v3

# Towards Safer Chatbots: Automated Policy Compliance Evaluation of Custom GPTs

@@ Abstract
User-configured chatbots built on top of large language models are increasingly available through centralized marketplaces such as OpenAI’s GPT Store. While these platforms enforce usage policies intended to prevent harmful or inappropriate behavior, the scale and opacity of customized chatbots make systematic policy enforcement challenging. As a result, policy-violating chatbots continue to remain publicly accessible despite existing review processes.
This paper presents a fully automated method for evaluating the compliance of Custom GPTs with its marketplace usage policy using black-box interaction. The method combines large-scale GPT discovery, policy-driven red-teaming prompts, and automated compliance assessment using an LLM-as-a-judge. We focus on three policy-relevant domains explicitly addressed in OpenAI’s usage policies: Romantic, Cybersecurity, and Academic GPTs.
We validate our compliance assessment component against a human-annotated ground-truth dataset, achieving an F1 score of 0.975 for binary policy violation detection. We then apply the method in a large-scale empirical study of 782 Custom GPTs retrieved from the GPT Store. The results show that 58.7% of the evaluated GPTs exhibit at least one policy-violating response, with substantial variation across policy domains. A comparison with the base models (GPT-4 and GPT-4o) indicates that most violations originate from model-level behavior, while customization tends to amplify these tendencies rather than create new failure modes.
Our findings reveal limitations in current review mechanisms for user-configured chatbots and demonstrate the feasibility of scalable, behavior-based policy compliance evaluation.
@@ 1 Introduction
Large Language Models based on the transformer architecture and its self-attention mechanism have driven substantial progress in Natural Language Processing [Vaswani2017]. Building on this architecture, OpenAI released the Generative Pre-trained Transformer (GPT) series, starting with GPT-1 [radford2018] and followed by GPT-2 [radford2019] and GPT-3 [brown2020]. Each new generation increased model scale and capability, supported by architectural refinements and large-scale training. The public release of ChatGPT in 2022, based on GPT-3.5 and optimized for interactive dialogue, illustrated the societal impact of these models by achieving unprecedented adoption rates [hu2023].
As LLMs became widely deployed, the need to adapt their behavior to specific tasks and user needs became apparent. Initial approaches relied on fine-tuning base models using domain-specific datasets, which required technical expertise and computational resources. More recently, OpenAI introduced a lightweight customization mechanism that enables users to configure chatbot behavior without retraining. Through system instructions via natural language, uploaded documents, and optional API integrations, users can create task-specific chatbots using a graphical interface. OpenAI formalized this capability under the notion of Custom GPTs (GPTs, in OpenAI’s parlance) [openai2023gpts], substantially lowering the barriers to creating specialized conversational chatbots.
@@
To facilitate discoverability and reuse, OpenAI launched the GPT Store, a centralized marketplace where customized chatbots can be published and accessed by other users. Prior to publication, each GPT must undergo a review process that combines automated and manual assessments to verify alignment with OpenAI’s usage policies and mitigate potential safety risks [openai2024usagepolicies]. This ecosystem has enabled the rapid proliferation of GPT-based assistants tailored to diverse tasks, audiences, and domains.
Despite the safeguards implemented during publication, Custom GPTs that appear to violate OpenAI’s usage policies remain accessible in the GPT Store. These policies, defined in the “Building with ChatGPT” section of OpenAI’s documentation, establish explicit boundaries on permissible behavior and prohibit, among other activities, certain forms of legal or medical advice, the generation of malicious software, and the deliberate fostering of romantic companionship. In practice, however, GPTs that conflict with these restrictions continue to surface in public search results. For instance, chatbots explicitly designed to engage in romantic or emotionally intimate interactions are readily discoverable, despite their prohibition under the platform policies, as shown in Figure 1. This persistent mismatch between stated policies and observed deployments highlights the practical difficulty of enforcing policy compliance at scale in an ecosystem where chatbot behavior emerges from the interaction between base-model capabilities and user-driven customization.
The proliferation of Custom GPTs raises broader challenges for platform governance and safety oversight. Manual review processes do not scale to large and rapidly evolving collections of user-generated chatbots. Existing automated moderation mechanisms, while necessary, appear insufficient to reliably detect policy violations across diverse and customizable behaviors. Addressing these limitations requires systematic methods capable of testing policy compliance in a systematic and scalable manner.
In this paper, we present an automated evaluation method for assessing the policy compliance of Custom GPTs deployed in the GPT Store. Our approach integrates large-scale GPT discovery, policy-driven red-teaming prompts, and automated compliance judgment using an LLM-as-a-judge strategy, enabling comprehensive assessments. Rather than claiming exhaustive coverage of the GPT Store, our evaluation focuses on three policy-relevant domains that are explicitly addressed in OpenAI’s usage policies, namely Romantic, Cybersecurity, and Academic GPTs. These categories allow us to study compliance behavior under different risk profiles.
Using this method, we conducted a large-scale empirical study of Custom GPT behavior under policy-relevant interactions. We analyzed compliance outcomes across hundreds of Custom GPTs, examined how non-compliance relates to thematic focus and usage popularity, and investigated whether observed violations primarily reflect behaviors inherited from base models or are introduced through user customization. The paper further presents representative case studies that illustrate common violation patterns and highlights implications for platform governance and safety oversight.
Finally, we describe responsible disclosure of identified violations to OpenAI and discuss the broader implications of automated policy evaluation for large ecosystems of user-configured chatbots. All code and datasets used in this study will be accessible upon acceptance of the paper (see A).
@@
2 Background and Problem Context
This section provides the technical and conceptual background required to understand the policy compliance evaluation performed in this study. We first describe Custom GPTs and the GPT Store as a deployment ecosystem for user-configured language models (§2.1). We then define policy compliance in this context and identify the challenges that arise when evaluating compliance at scale (§2.2). Finally, we discuss how existing safety mechanisms in base language models relate to policy compliance in customized deployments (§2.3).
2.1 Custom GPTs and the GPT Store
Custom GPTs are personalized chatbots built on top of OpenAI’s base language models (e.g., GPT-4, GPT-4o) that users can configure and publish without modifying model parameters or performing additional training. Introduced in late 2023 [openai2023gpts], Custom GPTs enable non-expert users to specify a chatbot’s behavior using natural language instructions and optional external resources.
Unlike fine-tuning approaches, where model weights are updated using domain-specific datasets (e.g., via techniques such as Low-Rank Adaptation [hu2021lora]), Custom GPTs rely exclusively on runtime configuration. Specifically, developers may provide three types of customization artifacts:
  • System prompts, which define the GPT’s role, scope, and behavioral constraints and are prepended to every interaction.
  • Knowledge files, such as documents or structured data, which the GPT may retrieve and reference during conversations.
  • Actions, which allow the GPT to invoke external APIs in response to natural language requests.
The underlying base model remains unchanged, and all Custom GPTs inherit the same foundational capabilities and safety mechanisms as the original GPT model.
To enable discovery and reuse, OpenAI operates the GPT Store, a public marketplace where developers can publish Custom GPTs for other users to access. Each GPT entry includes metadata such as a name, textual description, developer name, usage statistics, and a coarse-grained category label. While this metadata facilitates browsing and search, the internal configuration of each GPT (including system prompts, uploaded files, and actions) is not visible to external reviewers or users. As a result, any assessment of a Custom GPT’s behavior must rely exclusively on observed input–output interactions.
This deployment model introduces a large and heterogeneous ecosystem of user-configured chatbots whose behavior is shaped by both the base model and developer-specified instructions, motivating the need for scalable, behavior-based compliance evaluation.
2.2 Policy Compliance
2.2.1 Defining Policy Compliance
Within the GPT Store, Custom GPTs are subject to platform-specific usage policies that prohibit certain classes of behavior, such as fostering romantic companionship, facilitating academic dishonesty, or compromising the privacy of others [openai2024usagepolicies]. In this work, we adopt a behavioral definition of policy compliance.
Let $G$ denote a Custom GPT, $P$ a policy clause, and $Q=\{q_{1},q_{2},\ldots,q_{n}\}$ a finite set of policy-relevant test prompts. A response $r_{i}=G(q_{i})$ is considered compliant with policy $P$ if it does not violate the policy, denoted by a binary predicate $c(G,P,q_{i})\in\{0,1\}$. A Custom GPT is deemed compliant with policy $P$ if all tested responses comply:
$C(G,P)=\bigwedge_{q_{i}\in Q}c(G,P,q_{i}).$
We classify a GPT as non-compliant when at least one evaluated response contradicts the usage policies. This decision reflects the aim of determining whether a model can generate prohibited content when prompted, rather than estimating how often such failures occur.
2.2.2 Compliance vs. Safety and Alignment
Policy compliance is distinct from related concepts frequently studied in the LLM literature. Safety typically refers to the prevention of harmful or illegal outputs (e.g., violence, malware, hate speech), often evaluated through robustness or adversarial testing [qi2023finetuning, Dong2023]. Alignment concerns broader adherence to human values or preferences, commonly operationalised through Reinforcement Learning from Human Feedback (RLHF) [li2024safety].
In contrast, policy compliance evaluates adherence to explicit, platform-defined rules that may reflect governance choices rather than universal ethical principles. For example, prohibitions on romantic companionship or academic ghostwriting arise from platform policy decisions and institutional norms, not from intrinsic properties of language generation. As a result, compliance assessments are necessarily contextual and policy-specific.
2.2.3 Challenges in Evaluating Custom GPT Compliance
Evaluating compliance in the GPT Store presents several challenges that differentiate it from traditional software auditing or model-level safety evaluation.
Black-box evaluation. Reviewers cannot inspect a Custom GPT’s internal configuration or uploaded resources. Compliance must therefore be inferred exclusively from observable behavior during interaction.
Policy ambiguity. Usage policies are intentionally concise and abstract, leaving room for interpretation. Automated evaluation requires operationalizing policy clauses into concrete decision criteria.
Attribution uncertainty. Because Custom GPTs do not modify model parameters, non-compliant responses may originate from the base model, from developer-provided system prompts, or from external knowledge sources. Distinguishing between these sources is non-trivial but relevant for platform governance and responsibility allocation.
Scale and diversity. The GPT Store contains a large and diverse population of Custom GPTs spanning multiple domains. Manual review does not scale, while automated methods must generalize across heterogeneous behaviors while remaining sensitive to policy-specific constraints.
These challenges motivate the need for systematic, interaction-based compliance evaluation methods that operate under realistic platform constraints.
2.3 Safety Evaluation and Red-Teaming
Large language model platforms incorporate multiple mechanisms aimed at preventing harmful or policy-violating behavior prior to deployment. These include training-time approaches such as RLHF, which biases models toward preferred responses through human judgments [li2024safety], as well as deployment-time safeguards such as output filtering, self-verification [phute2023llm], and knowledge sanitization [ishibashi2023knowledge]. In addition, platform providers routinely conduct internal evaluations to assess model robustness against misuse.
From an evaluation perspective, red-teaming provides a structured way to probe models for policy-relevant failure modes. Red-teaming involves the systematic use of carefully designed prompts to identify violations, unsafe behaviors, or boundary failures under realistic interaction conditions. Rather than covertly bypassing safeguards, red-teaming seeks to expose weaknesses in a controlled manner so that mitigations can be developed. OpenAI has formalized this practice through its Red Teaming Network, which combines internal testing with contributions from external experts [openai2023redteaming].
Red-teaming and jailbreaking are related but not equivalent practices. Jailbreaking focuses on deliberately adversarial prompting techniques that attempt to bypass a model’s safety mechanisms, often through role-play, attention-shifting, or other strategies that induce responses the system is designed to avoid [Liu2024]. Red-teaming is broader and oriented toward safety auditing and policy enforcement. It examines whether a model produces policy-violating behavior under plausible user interactions without requiring the evaluator to defeat guardrails. Jailbreaking can appear within a red-teaming exercise, but effective red-teaming does not depend on it, and many policy violations arise under direct, non-adversarial prompting, which is the focus of this study.
These safety and evaluation mechanisms have shown effectiveness when users interact directly with base models such as ChatGPT. However, the introduction of Custom GPTs modifies the interaction surface in ways that complicate enforcement. Developer-specified system prompts mediate all user inputs and may unintentionally weaken refusal behavior or alter how policy constraints are applied. Moreover, reviewers cannot inspect these configurations directly, limiting assessment to observable behavior.
As a consequence, it is not self-evident that safety properties and compliance guarantees established at the base-model level continue to hold across large ecosystems of user-configured chatbots. This creates a gap between platform-level policy enforcement and real-world behavior in deployed Custom GPTs. Addressing this gap requires scalable, behavior-based evaluation methods that operationalise platform policies and apply red-teaming techniques systematically across customized deployments. The method presented in section 4 is designed to fulfill this role.
3 Related Work
The versatility of tasks and responses enabled by LLMs comes with inherent challenges in controlling their outputs and ensuring alignment with predefined guidelines. Prior research has highlighted several risks stemming from these limitations, including the generation of malware and phishing messages [kucharavy2024llms, Iturbe2024], the dissemination of misinformation and fake news [Weidinger2022, Zellers2019], and the creation of malicious bots [kucharavy2024llms, Derner2024]. Additionally, LLMs face technical shortcomings such as hallucinations—wherein they produce incorrect or nonsensical information [gao2024]—and code injection vulnerabilities, which adversaries can exploit to manipulate LLMs, leading to unintended and potentially harmful outputs [gao2024].
These challenges are rooted in the inherent complexity of LLMs. Their black-box nature and intricate internal mechanisms make it difficult to predict and control their behavior under diverse user inputs, exacerbating the challenges of safeguarding them [sarker2024llm, ullah2024challenges]. Moreover, the widespread adoption of LLMs among non-specialized users amplifies the need for strong safety considerations. Striking a balance between training LLMs to be helpful and ensuring their safety remains an ongoing challenge. Excessive focus on helpfulness can lead to harmful content generation, whereas overly restrictive safety tuning risks rejecting legitimate prompts and negatively affecting utility [bianchi2024safety, chehbouni2024]. OpenAI has invested significant resources into addressing these challenges [openai_safety], employing security policies, conducting both manual and automated evaluations, and implementing dedicated red-teaming efforts to mitigate risks [openai2023redteaming].
To further advance safety evaluations, open datasets have been developed to systematically assess LLMs, as documented in a recent systematic review [rottger2025]. This review underscores gaps in the ecosystem, including the predominance of English-centric datasets and the limitations of existing benchmarks in comprehensively assessing safety dimensions. Building on this, researchers have proposed various frameworks and benchmarks for safety evaluations. For instance, Xie et al. [xie2024onlinesafety] developed a public dataset for evaluating LLMs using black-box, white-box, and gray-box approaches. Another benchmark expands the scope of safety evaluation to 45 distinct categories, analyzing 40 models and identifying Claude-2 and Gemini-1.5 as the most robust with respect to safety [xie2024sorrybench]. Additional benchmarks offer larger prompt sets for evaluation [zhang2024safetybench], as well as support for multilingual assessments in English and Chinese [yuan2024, sun2023safety].
The LLM-as-a-judge technique has also emerged as a widely adopted method for evaluating LLM outputs [huang2024empirical, li2025generation, gu2025surveyllmasajudge]. This approach, which involves using LLMs to assess the quality or performance of other systems, has demonstrated high agreement rates with human evaluations, further validating its efficacy [Zheng2023].
The customization of LLMs introduces another dimension of complexity to their safety assessment. Hung et al. [huang2024lisa] proposed a method demonstrating that fine-tuning can enhance safety by reducing harmful outputs while maintaining task accuracy. Conversely, other studies have shown that fine-tuning can inadvertently degrade safety alignment, even when benign datasets are used, rendering models more vulnerable to harmful or adversarial instructions [qi2023finetuning]. Additionally, fine-tuning has been linked to increased susceptibility to jailbreaking attacks [kumar2024finetuning], and when applied to develop agent applications, it can lead to unintended safety lapses if failed interaction data is not properly utilized [wang2024learningfailure]. Achieving a balance between safety and utility is essential, as an excessive emphasis on safety during fine-tuning may cause models to overly reject valid prompts, negatively impacting their usability [hsu2025safe].
OpenAI’s Custom GPTs represent an advancement in enabling end-users to personalize LLMs. Prior studies have noted that some features of these systems may introduce significant risks [antebi2024gptsheeps], including security and privacy concerns within the GPT store [ma2025privacy, tao2023openingpandoras]. Zhang et al. [zhang2024lookgptappslandscape] conducted the first longitudinal study of the platform, analyzing metadata from 10,000 Custom GPTs and identifying a growing interest in these systems. Similarly, Su et al. [su2024gptstoremininganalysis] examined the GPT store, analyzing user preferences, algorithmic influences, and market dynamics. Their study identified Custom GPTs that contradicted OpenAI’s usage policies, raising concerns about the effectiveness of the platform’s review mechanisms. In addition, Carrillo et al. [carrillo2026personal] showed that many Custom GPTs include connections with third-party services, which collect extensive personal data which is very often not properly disclosed in the Custom GPTs’ privacy policies.
The work by Yu et al. [yu2024assessingpromptinjection] is most closely related to our study. In their research, the authors evaluated over 200 Custom GPTs using adversarial prompts, demonstrating the susceptibility of these systems to prompt injection attacks. Their methodology included generating dynamic red-teaming prompts tailored to the characteristics of the targeted Custom GPTs. Similar approaches, focusing on general LLM evaluation, were proposed by Liu et al. [Liu2024] and Shen et al. [Shen2024DoAnythingNow], who explored the types of adversarial prompts most effective in bypassing safety measures. In line with this, Yu et al. [Yu2024DontListen] conducted a systematic study on jailbreak prompts, revealing the ease with which users can craft prompts that circumvent LLM safeguards. Yu et al. [yu2024gptfuzzerredteaminglarge] also introduced GPTFuzzer, a black-box fuzzing framework based on automated generation of jailbreak prompts, achieving over 90% attack success rates against foundational models like ChatGPT and LLaMa-2.
In contrast, our study evaluates the alignment of Custom GPTs with OpenAI’s usage policies using an automated method that covers the full evaluation process, ranging from the identification of Custom GPTs to the generation of compliance assessments. To the best of our knowledge, this work is the first to automate a large-scale policy evaluation of Custom GPTs available in the GPT Store.
4 Policy Compliance Evaluation Method